OpenStudy (anonymous):

S=16Ad^-2/3 Write d in terms of A and S. Anyone know how ?

jhonyy9 (jhonyy9):

so first divide boths sides by 16A and will get S/16A = 1/cuberoot(d^2) so because d^-2/3 = 1/cuberoot(d)^2 cross multiplie and will get S*cuberoot(d^2) =16A than now divide both sides by S cuberoot(d^2)=16A/S make on exponent 3 both sides d^2 = (16A/S)^3 make squareroot from both sides d=(+/-) sqrt((16A/S)^3) = 16A/S *(+/-)sqrt(16A/S) = 4*16A/S *(+/-)sqrt(A/S) = =64A/S *(+/-)sqrt(A/S) so this (+/-) sign that like for example sqrt(9)= +/- 3 so because 3*3=9 and (-3)*(-3)=9 too hope so much that is understandably sure right good luck bye